OpenStudy (loser66):

which amount of money can be formed from 2dollar bill and 5 dollar bill? Please, help

OpenStudy (just_one_last_goodbye):

Um... $10

OpenStudy (just_one_last_goodbye):

$20

OpenStudy (salty):

2n+5m n,m->I

OpenStudy (loser66):

Thank you, but I need to know the formula of n, where n = 2x+5y

OpenStudy (salty):

what do you mean by formula for n? isn't it just n=2x+5y where x and y are integers

OpenStudy (marie152347xoxo):

20

OpenStudy (just_one_last_goodbye):

\[x = \frac{ n - 5y }{ 2 }\]

OpenStudy (loser66):

2x+ 5y = n So, the constraint is P(4) , since if n=4, then x =2,y =0

OpenStudy (loser66):

for n =5, then x = 0, y =1

OpenStudy (loser66):

But if n =3, you don't have P(3) because there is no x,y satisfy 2x+5y =3

OpenStudy (loser66):

P(4) and up checked Now , by strong induction, we need to find the formula of n.

OpenStudy (loser66):

\(P(4)\cap P(5)\cap P(6)\cap \cdots \cap P(n)\rightarrow P(n+1)\)

OpenStudy (loser66):

\(P(n-3)P(n-2)P(n-1) = P(n) - stuck!!) = P(n+1)\)

OpenStudy (loser66):

And I need help to unlock the point I stuck. Just from there, I can construct the formula of n and prove it by induction.

OpenStudy (loser66):

I got it

OpenStudy (loser66):

Let P(n) := we can form all amounts of money greater or equal to 4 from 2 dollar bill and 5 dollar bill P(4) = 2*(2) +0*(5) P(5) =0*(2) +1*(5) So base cases are true.

OpenStudy (loser66):

Hypothesis step: Suppose P(k) is true for k amount of money so, we need prove P(k+1) From HS, \(4\leq j\leq k\) So P(k-1) is true. if P(k-1) is true, now we add one 2 dollar bill in that amount, we have k-1+2 = k+1, so the amount of k+1 is formed by 2 dollar bill and 5 dollar bill also. Proof done.